A 15.00 g sample containing mixed alkali and other inert components was dissolved and diluted to 300 mL with water. A 20 mL aliquot was titrated with 5.02 mL of 0.5352 M HCl to reach PHP endpoint. Another 20 mL aliquot was titrated to the BCG endpoint, using up 18.87 mL of titrant in the process. How much volume of titrant was needed to reach the second endpoint starting from the first endpoint? Answer in two decimal places.


1
Expert's answer
2021-05-03T07:56:28-0400

V = "\Delta"V = V2 - V1 = 18.87 - 5.02 = 13.81 ml.

Answer: V = 13.81 ml.